Starting on October 1, 2003, annual deposits of $155 are made into an account paying interest at a rate of 8.4% compounded monthly. How much is in the account immediately after the deposit on October 1, 2030? *do not round intermediate steps*
Answer #1

Number of deposits = 2030 - 2003 + 1 = 28

n = 28

r = (1 + 0.084/12)^12 - 1 = 1.0873106619 - 1

r = 0.0873106619

r = 8.73106619%

PMT = 155

FV = \frac{PMT}{r} * \left [ (1+ r )^{n} - 1 \right ]

FV = \frac{155}{0.0873106619} * \left [ (1+ 0.0873106619 )^{28} - 1 \right ]

FV = 1,775.2700143028* 9.4208250749

FV = 16,724.5082654619

Can you please upvote? Thank You :-)